# Applying the Liouville-Lanczos Method of Time-Dependent Density-Functional Theory to Warm Dense Matter

### v1

@@ II.1 Dynamical structure factor and imaginary-time density–density correlation function
The main computational burden of LR-TDDFT lies in solving Dyson’s equation for the density response function. The standard way is to compute $\chi^{0}_{\scriptscriptstyle\mathbf{G}\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ and use matrix inversion to solve Dyson’s equation (2). An alternative approach is the LL method, which finds the solution by employing an iterative approach derived from time-dependent density-functional perturbation theory using the quantum Liouville equation for the reduced one-electron KS density matrix (see Refs. Rocca et al. (2008); Timrov et al. (2015); Ge et al. (2014); Saad (2003)). Theoretically, these approaches should yield the same solution for $\chi(\mathbf{q},\omega)$. However, variations in implementation details—such as the choice of basis sets for wavefunctions, handling of tightly bound electrons, and the numerical solvers used—can introduce numerical differences that lead to deviations between the results. As far as we are aware, the LL method for $\chi(\mathbf{q},\omega)$ has only been implemented in Quantum ESPRESSO Timrov et al. (2015), which utilizes pseudopotential-based KS-DFT with normconserving and ultrasoft pseudopotentials. This implementation is focused on the applications at ambient conditions to model the energy loss spectrum of electrons in solids. To test this implementation of the LL method at WDM conditions, we compare its results for the DSF with the data computed using the standard LR-TDDFT based on the PAW formulation. Furthermore, we validate the LL approach by examining the ITCF results and comparing them with the data from PIMC simulations for warm dense hydrogen.
@@
As mentioned above, LR-TDDFT does not capture the quasi-inelastic part of the DSF, which translates to a constant shift of the ITCF that does not depend on $\tau$. To circumvent this aspect, we compare the shifted ITCF values:
@@
In practice, the Lorentzian smearing parameter $\eta$ is introduced into the LR-TDDFT calculations to regularize $\chi(\mathbf{q},\omega)$ (both in the standard LR-TDDFT and the LL method). This is done by adding a small imaginary part to the frequency, transforming it from $\omega$ to $\omega+i\eta$. As a result, discrete spectral features (lines) in the DSF are broadened Rocca et al. (2008). In the real frequency domain, this technique serves as a method to compute a continuous $\chi(\mathbf{q},\omega)$, with a negligible impact on the shape and features of the DSF $S(\mathbf{q},\omega)$. However, the effect of Lorentzian smearing on the ITCF $F(\mathbf{q},\tau)$, computed using $S(\mathbf{q},\omega)$ from LR-TDDFT, has not been considered in previous studies. This is a relevant question because XRTS measurements can be analyzed in the Laplace domain, which helps to circumvent the deconvolution problem of the XRTS spectrum with the source and instrument function $R(\omega)$ Dornheim et al. (2023b, e). Therefore, we also examine the impact of Lorentzian smearing employed in LR-TDDFT on the ITCF.
@@ III.1 Isochorically heated aluminum
We begin by analyzing the DSF of isochorically heated Al, which has cold ions arranged in the fcc lattice structure and hot electrons at a temperature of $T=6~{}{\rm eV}$. Such a transient state is generated and examined in experiments that utilize X-ray-driven heating used for XRTS measurements Descamps et al. (2024, 2020); Nicoul et al. (2011); White et al. (2014). In these experiments, the incident X-rays do not directly interact with the ions, keeping them cold during a free electron laser pulse that lasts less than 100 femtoseconds.
@@
Figure 4: The top panel shows the density distribution in the simulation cell with $N=14$ protons at $r_{s}=2$ and $T=12.58~{}{\rm eV}$. The middle panel illustrates the convergence of the DSF calculations of warm dense hydrogen with respect to the number of Lanczos iterations in the LL method. The bottom panel compares the DSF results computed using ultrasoft (denoted as US) and norm-conserving (denoted as NC) pseudopotentials in the LL method with standard LR-TDDFT within the PAW framework.
In Fig. 1, we show the convergence of the results for isochorically heated Al with respect to the number of iterations $N_{\rm iter}$ in the LL calculations at $q=1.396~{}{\rm\AA^{-1}}$. From Fig. 1, it is evident that the LL results are converged after $N_{\rm iter}=2000$ iterations. In Fig. 2, considering the LDA level description of the XC functional, the LL results computed using $N_{\rm iter}=3000$ are compared with the exact solution of Dyson’s equation within the PAW formalism. Overall, we find good agreement between the different approaches. At $q=0.155~{}{\rm\AA^{-1}}$, one can see a well-defined plasmon feature, which becomes broader with the increase in the wavenumber, indicating a stronger Landau damping effect Landau (1946). The pair continuum for Al corresponds to $q>1.3~{}{\rm\AA^{-1}}$, where the plasmon is not well defined anymore Giuliani and Vignale (2008); Hamann et al. (2020). A small spike at $\omega<10~{}{\rm eV}$ at $q=0.7757~{}{\rm\AA^{-1}}$ and $q=1.396~{}{\rm\AA^{-1}}$ is due to lattice effects leading to energy gaps at Brillouin zone boundaries Larson et al. (2000). An interesting observation from Fig. 2 is a noticeable difference in the magnitude of the DSF maximum for $q=0.7757~{}{\rm\AA^{-1}}$ between the LL method and Dyson’s equation solution within the PAW formalism. To analyze this aspect further, we compare the results of the DSF computed using the PBE XC functional with those obtained using LDA level functionals, as shown in Fig. 3. The calculations were performed using norm conserving pseudopotentials and PAW setups that are consistent with the XC functionals used. The PBE-based results from the LL method are in close agreement with the LDA-level calculations using the same method. When examining the DSFs obtained from solving Dyson’s equation within the PAW framework, we find that the maximum of the DSFs calculated with the PBE functional are larger than those derived from the LDA-level data. This larger value leads to closer agreement with the results obtained from the LL method. We conclude that the observed differences in the vicinity of the DSF maximum are due to the differing methods (pseudopotentials) used to handle core electrons.
@@ III.2 Warm dense hydrogen at metallic density
Figure 6: Top panel: DSF of warm dense hydrogen computed using 10 different snapshots of $N=14$ protons (red lines) and corresponding averaged result (green line). Bottom panel: the shifted ITCF results defined by Eq. (4) from the calculations using the LL approach, LR-TDDFT and PIMC. The DSF and ITCF results are compared with the data computed for a UEG model Dornheim et al. (2020). The LR-TDDFT results were computed at $r_{s}=2$, $T=12.58~{}{\rm eV}$, and using the Lorentzian smearing parameter $\eta=0.2~{}{\rm eV}$.
@@
Figure 8: Top panel: DSF of warm dense hydrogen computed using 10 different snapshots of $N=14$ protons (red lines) and corresponding averaged result (green line) at $\eta=2~{}{\rm eV}$. Middle panel: effect of the variation of the Lorentzian smearing parameter in the range $0.1~{}{\rm eV}\leq\eta\leq 3.0~{}{\rm eV}$. Bottom panel: shifted ITCF results defined by Eq. (4) from calculations using the LL approach, LR-TDDFT (with different $\eta$ values) and PIMC. The DSF and ITCF results are compared with the data computed using the UEG model Dornheim et al. (2020). The results are presented for $r_{s}=2$ and $T=12.58~{}{\rm eV}$.
@@
In the middle panel of Fig. 8, we present the results for the snapshot-averaged DSF at $q=4.584~{}\text{Å}^{-1}$ for $0.1~{}\text{eV}\leq\eta\leq 3.0~{}\text{eV}$. Notably, we observe significantly stronger quasi-periodic noise in the DSF for $\eta<1~{}\text{eV}$ compared to that at $q=1.528~{}\text{Å}^{-1}$ (see Fig. 7). A smooth curve for the DSF is produced using $\eta\gtrsim 2~{}\text{eV}$. We note that the magnitude of $\eta$ should be gauged taking into account the characteristic frequency (energy) scale $\omega$ of the DSF. For example, the full width at half maximum (FWHM), denoted as $\Delta\omega_{0}$, at $q=4.584~{}\text{Å}^{-1}$ is approximately 3.5 times larger than that at $q=1.528~{}\text{Å}^{-1}$. The relative value of $\eta/\Delta\omega$ required for damping the noise in the DSF is similar for both wavenumbers, with an optimal value around $\eta_{0}=10^{-2}\times\Delta\omega_{0}$. This smearing degree results in a minor effect on the overall shape of the DSF, with the magnitude of the DSF maximum being underestimated by about $5\%$. The FWHM $\Delta\omega_{0}$ and the position of the DSF maxima, $\omega_{0}$, remain nearly unchanged for both $q=4.584~{}\text{Å}^{-1}$ and $q=1.528~{}\text{Å}^{-1}$.
@@ IV Conclusions
To test the LL method, we first considered isochorically heated Al, which features hot electrons in an fcc lattice structure of cold ions. Such transient states are generated and examined in experiments that utilize X-ray-driven heating Descamps et al. (2024, 2020); Nicoul et al. (2011); White et al. (2014) (also see Ref. Moldabekov et al. (2024a) for detailed discussions). For the fcc Al, we found good overall agreement between the results from the LL method and the solution of Dyson’s equation in the standard LR-TDDFT within the PAW formalism. However, we noted differences in the DSF values near the DSF maximum between the two methods. This discrepancy arises from the different approaches in approximating the interaction of core electrons with valence electrons. Specifically, the pseudopotential or PAW setup employed defines the interaction of valence electrons with the crystal lattice, affecting the band structure and, consequently, the DSF as well as EELS Smith and Segall (1986); Schülke et al. (1995, 1993); Weissker et al. (2010); Moldabekov et al. (2024a). Our analysis of the DSF for fcc Al demonstrated that if relatively subtle features of the DSF are of interest, particular care must be taken by cross-validating results using different approaches to TDDFT calculations.

### v2

@@ II.1 Dynamical structure factor and imaginary-time density–density correlation function
The main computational burden of LR-TDDFT lies in solving Dyson’s equation for the density response function. The standard way is to compute $\chi^{0}_{\scriptscriptstyle\mathbf{G}\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ and use matrix inversion to solve Dyson’s equation (2). An alternative approach is the LL method, which finds the solution by employing an iterative approach derived from time-dependent density-functional perturbation theory using the quantum Liouville equation for the reduced one-electron KS density matrix (see Refs. Rocca et al. (2008); Timrov et al. (2015); Ge et al. (2014)).
The LL approach to LR-TDDFT utilizes the definition of the density response function in terms of the reduced one-electron KS density matrix Timrov et al. (2015):
$\chi(\mathbf{q},\omega)={\rm Tr}\Big{(}\hat{n}_{\mathbf{q}}(\omega)\delta\hat{\rho}_{\mathbf{q}}(\omega)\Big{)},$ (3)
where $\delta\hat{\rho}_{\mathbf{q}}(\omega)=\hat{\rho}_{\mathbf{q}}(\omega)-\hat{\rho}^{0}$ denotes the perturbation in the KS density matrix due to a weak perturbation $\delta v_{\rm ext,\mathbf{q}}\sim e^{i\mathbf{q}\mathbf{r}}$ applied to an equilibrium state (defined by the unperturbed density matrix $\rho^{0}(\mathbf{r},\mathbf{r}^{\prime})$ and the KS Hamiltonian $\hat{H}^{0}_{\rm KS}$), and $\hat{n}_{\mathbf{q}}$ denotes the response of the $\mathbf{q}$ Fourier component of the charge-density operator.
To compute $\chi(\mathbf{q},\omega)$, instead of the time-dependent KS equation, a quantum Liouville equation for the reduced one-electron KS density matrix is considered in the LL method:
$i\frac{\mathrm{d}\hat{\rho}(t)}{\mathrm{d}t}=\left[\hat{H}_{\rm KS}(t),\hat{\rho}(t)\right],$ (4)
where the KS Hamiltonian $\hat{H}_{\rm KS}=-\frac{\nabla^{2}}{2}+v_{\rm ext}(\mathbf{r},t)+v_{\rm HXC}(\mathbf{r},t)$ contains the total external potential, $v_{\rm ext}(\mathbf{r},t)$, due to the ions and any additional perturbations, and $v_{\rm HXC}(\mathbf{r},t)$ corresponding to the Hartree and exchange-correlation contributions. The KS density matrix is represented by the KS orbitals $\phi_{i}$ and corresponding occupation numbers $f_{i}$ as $\rho(\mathbf{r},\mathbf{r}^{\prime},t)=\sum_{i}f_{i}\phi_{i}(\mathbf{r},t)\phi_{i}(\mathbf{r}^{\prime},t)$.
Considering a weak perturbation $\delta v_{\rm ext}(\mathbf{q},\omega)$, linearization of Eq. (4) and subsequent Fourier transformation allows one to find:
$\left(\omega-\hat{L}\right)\cdot\delta\hat{\rho}_{\mathbf{q}}(\omega)=\left[\delta v_{\rm ext,\mathbf{q}}(\omega),\hat{\rho}^{0}\right],$ (5)
where the Liouvillian superoperator $\hat{L}$ is defined as
$\hat{L}\cdot\delta\hat{\rho}_{\mathbf{q}}(\omega)=\left[\hat{H}^{0}_{\rm KS},\delta\hat{\rho}_{\mathbf{q}}(\omega)\right]+\left[\delta v_{\rm HXC,\mathbf{q}}(\omega),\hat{\rho}^{0}\right].$ (6)
From Eq. (3), we see that $\chi(\mathbf{q},\omega)$ can be computed if one has the solution of Eq. (5), which can be written as
$\delta\hat{\rho}_{\mathbf{q}}(\omega)=\left(\omega-\hat{L}\right)^{-1}\cdot\left[\delta v_{\rm ext,\mathbf{q}}(\omega),\hat{\rho}^{0}\right]$ (7)
The LL approach is an iterative method computing a tridiagonal form of the matrix $L\simeq L^{(M)}$ (defining $\hat{L}$ in a given representation) in terms of the Lanczos coefficients (wit $M$ being the number of Lanczos iterations). More details of the Lanczos algorithm can be found, e.g., in Ref. Saad (2003), and specifics of the implementation for TDDFT (the Liouville-Lanczos method) are presented in Refs. Rocca et al. (2008); Malcıoğlu et al. (2011); Ge et al. (2014); Timrov et al. (2015, 2013).
In Dyson’s equation (2), calculating the response function $\chi^{~{}0}_{\scriptscriptstyle\mathbf{G},\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ requires matrix elements that represent transitions between states with a momentum difference significantly larger than $\mathbf{q}$ and an energy eigenvalue difference that is considerably greater than $\omega$ Yan et al. (2011). This necessity means that a substantial number of additional bands are needed beyond those required to compute the equilibrium (unperturbed) density matrix $\rho^{0}$. In contrast, the Lanczos algorithm used in the LL method only relies on information in $\rho^{0}$ and does not require any extra empty bands.
@@
Theoretically, the standard LR-TDDFT method solving Dyson’s equation and LL approach should yield the same solution for $\chi(\mathbf{q},\omega)$. However, variations in implementation details—such as the choice of basis sets for wavefunctions, handling of tightly bound electrons, and the numerical solvers used—can introduce numerical differences that lead to deviations between the results. As far as we are aware, the LL method for $\chi(\mathbf{q},\omega)$ has only been implemented in Quantum ESPRESSO Timrov et al. (2015), which utilizes pseudopotential-based KS-DFT with normconserving and ultrasoft pseudopotentials. This implementation is focused on the applications at ambient conditions to model the energy loss spectrum of electrons in solids. To test this implementation of the LL method at WDM conditions, we compare its results for the DSF with the data computed using the standard LR-TDDFT based on the PAW formulation. Furthermore, we validate the LL approach by examining the ITCF results and comparing them with the data from PIMC simulations for warm dense hydrogen.
@@
As mentioned above, LR-TDDFT does not capture the quasi-elastic part of the DSF, which translates to a constant shift of the ITCF that does not depend on $\tau$. To circumvent this aspect, we compare the shifted ITCF values:
@@
In practice, the Lorentzian smearing parameter $\eta$ is introduced into the LR-TDDFT calculations to regularize $\chi(\mathbf{q},\omega)$ (both in the standard LR-TDDFT and the LL method). This is done by adding a small imaginary part to the frequency, transforming it from $\omega$ to $\omega+i\eta$. As a result, discrete spectral features (lines) in the DSF are broadened Rocca et al. (2008). In the real frequency domain, this technique serves as a method to compute a continuous $\chi(\mathbf{q},\omega)$, with a negligible impact on the shape and features of the DSF $S(\mathbf{q},\omega)$. However, the effect of Lorentzian smearing on the ITCF $F(\mathbf{q},\tau)$, computed using $S(\mathbf{q},\omega)$ from LR-TDDFT, has not been considered in previous studies. This is a relevant question because XRTS measurements can be analyzed in the Laplace domain, which helps to circumvent the deconvolution problem of the XRTS spectrum with the source and instrument function $R(\omega)$ Dornheim et al. (2023b, e). Therefore, we also examine the impact of Lorentzian smearing employed in LR-TDDFT on the ITCF. We note that the broadening parameter $\eta$ is a feature of LR-TDDFT and is not needed in the PIMC simulations of the ITCF.
@@ III.1 Isochorically heated aluminum
We begin by analyzing the DSF of isochorically heated Al, which has cold ions arranged in the fcc lattice structure and hot electrons at a temperature of $T=6~{}{\rm eV}$. Such a transient state is generated and examined in experiments that utilize X-ray-driven heating used for XRTS measurements Descamps et al. (2024, 2020); Nicoul et al. (2011); White et al. (2014). In these experiments, the incident X-rays do not directly interact with the nuclei, keeping them cold during a free electron laser pulse that lasts less than 100 femtoseconds.
@@
Figure 4: The top panel shows the density distribution in the simulation cell (in units of mean density $n_{0}$) with $N=14$ protons at $r_{s}=2$ and $T=12.58~{}{\rm eV}$. The middle panel illustrates the convergence of the DSF calculations of warm dense hydrogen with respect to the number of Lanczos iterations in the LL method. The bottom panel compares the DSF results computed using ultrasoft (denoted as US) and norm-conserving (denoted as NC) pseudopotentials in the LL method with standard LR-TDDFT within the PAW framework.
In Fig. 1, we show the convergence of the results for isochorically heated Al with respect to the number of iterations $N_{\rm iter}$ in the LL calculations at $q=1.396~{}{\rm\AA^{-1}}$. From Fig. 1, it is evident that the LL results are converged after $N_{\rm iter}=2000$ iterations. In Fig. 2, considering the LDA level description of the XC functional, the LL results computed using $N_{\rm iter}=3000$ are compared with the exact solution of Dyson’s equation within the PAW formalism. Overall, we find good agreement between the different approaches. At $q=0.155~{}{\rm\AA^{-1}}$, one can see a well-defined plasmon feature, which becomes broader with the increase in the wavenumber, indicating a stronger Landau damping effect Landau (1946). The pair continuum (strong plasmon damping) for Al corresponds to $q>1.3~{}{\rm\AA^{-1}}$, where the plasmon is not well defined anymore Giuliani and Vignale (2008); Hamann et al. (2020). A small spike at $\omega<10~{}{\rm eV}$ at $q=0.7757~{}{\rm\AA^{-1}}$ and $q=1.396~{}{\rm\AA^{-1}}$ is due to lattice effects leading to energy gaps at Brillouin zone boundaries Larson et al. (2000). An interesting observation from Fig. 2 is a noticeable difference in the magnitude of the DSF maximum for $q=0.7757~{}{\rm\AA^{-1}}$ between the LL method and Dyson’s equation solution within the PAW formalism. To analyze this aspect further, we compare the results of the DSF computed using the PBE XC functional with those obtained using LDA level functionals, as shown in Fig. 3. The calculations were performed using norm conserving pseudopotentials and PAW setups that are consistent with the XC functionals used. The PBE-based results from the LL method are in close agreement with the LDA-level calculations using the same method. When examining the DSFs obtained from solving Dyson’s equation within the PAW framework, we find that the maximum of the DSFs calculated with the PBE functional are larger than those derived from the LDA-level data. This larger value leads to closer agreement with the results obtained from the LL method. We conclude that the observed differences in the vicinity of the DSF maximum are due to the differing methods (pseudopotentials) used to handle core electrons.
@@ III.2 Warm dense hydrogen at metallic density
Figure 6: Top panel: DSF of warm dense hydrogen computed using 10 different snapshots of $N=14$ protons (red lines) and corresponding averaged result (green line). Bottom panel: the shifted ITCF results defined by Eq. (9) from the calculations using the LL method and PIMC. The DSF and ITCF results are compared with the data computed for a UEG model Dornheim et al. (2020). The LR-TDDFT and PIMC results were computed at $r_{s}=2$, $T=12.58~{}{\rm eV}$, $q\simeq 1.528~{}{\rm\AA^{-1}}$, and using the Lorentzian smearing parameter $\eta=0.2~{}{\rm eV}$.
@@
Figure 8: Top panel: DSF of warm dense hydrogen computed using 10 different snapshots of $N=14$ protons (red lines) and corresponding averaged result (green line) at $\eta=2~{}{\rm eV}$. Middle panel: effect of the variation of the Lorentzian smearing parameter in the range $0.1~{}{\rm eV}\leq\eta\leq 3.0~{}{\rm eV}$. Bottom panel: shifted ITCF results defined by Eq. (9) from calculations using the LL method (with different $\eta$ values) and PIMC. The DSF and ITCF results are compared with the data computed using the UEG model Dornheim et al. (2020). The results are presented for $r_{s}=2$ and $T=12.58~{}{\rm eV}$.
@@
In the middle panel of Fig. 8, we present the results for the snapshot-averaged DSF at $q=4.584~{}\text{Å}^{-1}$ for $0.1~{}\text{eV}\leq\eta\leq 3.0~{}\text{eV}$. Notably, we observe significantly stronger quasi-periodic noise in the DSF for $\eta<1~{}\text{eV}$ compared to that at $q=1.528~{}\text{Å}^{-1}$ (see Fig. 7). A smooth curve for the DSF is produced using $\eta\gtrsim 2~{}\text{eV}$. We note that the magnitude of $\eta$ should be gauged taking into account the characteristic frequency (energy) scale $\omega$ of the DSF. For example, the full width at half maximum (FWHM), denoted as $\Delta\omega_{0}$, at $q=4.584~{}\text{Å}^{-1}$ is approximately 3.5 times larger than that at $q=1.528~{}\text{Å}^{-1}$. The relative value of $\eta/\Delta\omega$ required for damping the noise in the DSF is similar for both wavenumbers, with an optimal value around $\eta_{0}=10^{-2}\times\Delta\omega_{0}$. This smearing degree results in a minor effect on the overall shape of the DSF, with the magnitude of the DSF maximum being underestimated by about $5\%$. At $\eta\lesssim\eta_{0}$, the FWHM $\Delta\omega_{0}$ and the position of the DSF maxima, $\omega_{0}$, remain nearly unchanged for both $q=4.584~{}\text{Å}^{-1}$ and $q=1.528~{}\text{Å}^{-1}$. The increase in $\eta$ gradually deteriorates the DSF maximum prediction due to the DSF asymmetry relative to $\omega_{0}$.
@@ IV Conclusions
To test the LL method, we first considered isochorically heated Al, which features hot electrons in an fcc lattice structure of cold ions. Such transient states are generated and examined in experiments that utilize X-ray-driven heating Descamps et al. (2024, 2020); Nicoul et al. (2011); White et al. (2014) (also see Ref. Moldabekov et al. (2024a) for detailed discussions). For the fcc Al, we found good overall agreement between the results from the LL method and the solution of Dyson’s equation in the standard LR-TDDFT within the PAW formalism. However, we noted differences in the DSF values near the DSF maximum between the two methods. This discrepancy arises from the different approaches in approximating the interaction of core electrons with valence electrons. Specifically, the pseudopotential or PAW setup employed defines the interaction of valence electrons with the crystal lattice, affecting the band structure and, consequently, the DSF as well as EELS Smith and Segall (1986); Schülke et al. (1995, 1993); Weissker et al. (2010); Moldabekov et al. (2024a). Our analysis of the DSF for fcc Al demonstrated that if relatively subtle features of the DSF are of interest, particular care must be taken by cross-validating results using different approaches to TDDFT calculations. This was also pointed out in a recent work by Hentschel et al. Hentschel et al. (2025) considering pseudopotentials with 3 and 11 valence electrons.
@@
The primary computational limitation of the standard LR-TDDFT arises from the memory requirements to store matrices on simulation cores Enkovaara (2012). This can reach several terabytes when using around $10^{3}$ bands. When a relatively small number of bands is required, e.g., for wavenumbers $q\lesssim q_{q_{F}}$ and small-sized systems (typically containing about $N\sim 10$ atoms), the standard LR-TDDFT approach can be up to two orders of magnitude more efficient than the LL method. However, in cases involving a large number of bands or atoms, as discussed in this work, employing standard LR-TDDFT becomes computationally unfeasible or highly impractical. In contrast, the LL method has lower memory requirements and does not necessitate empty states, allowing efficient parallelization over thousands of cores. This makes it capable of handling larger wavenumbers, broader frequency ranges, and larger systems. A detailed discussion of the comparison of the computational costs between the LL method and LR-TDDFT, using bulk diamond as an example, can be found in Ref. Timrov et al. (2015).